OpenStudy (anonymous):

M(-1,6) is the midpoint of RS. If S has the coordinates (3,13), find the coordinates of R. Multiple choice: 1) (7,-1) 2) (-5,20) 3) (7,20) 4) (-5,-1)

OpenStudy (anonymous):

( Sx + Rx ) / 2 = Mx (Sy+Ry)/2 = My Mx = -1 Sx + Rx = -2 Sx = 3 3 +Rx = -2 Rx = -5 My = 6 Sy+Ry = 12 Sy = 13 13 + Ry = 12 Ry = -1 R(-5,-1)

OpenStudy (anonymous):

why does Sx + Rx = -2?

OpenStudy (anonymous):

Ya i didnt understand that either

OpenStudy (anonymous):

(sx + rx) / 2 = Mx (sx + rx) / 2 = -1 sx+rx = -2

OpenStudy (anonymous):

oh ok. i got it. (Sx+Rx)/2 = -1